OpenStudy (anonymous):

3 7/8+ 2 1/2 with the steps so i can get it please

OpenStudy (anonymous):

\[\left(3+\frac{7}{8}\right)+\left(2+\frac{1}{2}\right)=\frac{31}{8}+\frac{5}{2}=\frac{51}{8}=6\frac{3}{8} \]